# A stochastic volatility approximation for a tick-by-tick price model with mean-field interaction

### v2

@@ 3.3 Convergence of the joint price-volatility process
The representations (66), (67) and (68) imply that the pair $(Z_{N},Y_{N})$ is not a Markov process. To see this we observe that (20), with $\lambda^{i,\pm}$ in place of $\lambda^{\pm}$ implies that the $0,+\infty)^{N}-$valued process $\mathbf{X}^{\pm}(t)=(X_{i}^{\pm}(t))_{i=1}^{N}$ is a Markov process whose infinitesima generator $\mathcal{L}_{N}$ acts on a smooth $F:[0,+\infty)^{N}\rightarrow{\mathbb{R}}$ as follows: for $\mathbf{x}=(x_{1}^{\pm},\ldots,x_{N}^{\pm})$
@@
Note that if $(Z_{N}(t),Y_{N}(t))$ were a Markov process, then $m_{N}^{i,\pm}(\mathbf{X(t)})$ would also be a Markov process, since they are linked by an invertible linear transformation (except in the special case $\beta=\gamma=1$; but in this case $Z_{N}=0$ and the argument can be adapted by replacing the pair $m_{N}^{i,\pm}(\mathbf{X(t)})$ with the sum of the two components). This would imply that, for a smooth $\Psi:[0,+\infty)^{2}\rightarrow{\mathbb{R}}$,
$\mathcal{L}_{N}\Psi(m_{N}^{\pm}(\mathbf{x}))$
would be a function of $m_{N}^{\pm}(\mathbf{x})$. This is not the case. Indeed observe that, by permutation invariance, $T_{i}^{s}\Psi-\Psi$ does not depend on $i$, and therefore in $\mathcal{L}_{N}\Psi(m_{N}^{\pm}(\mathbf{x}))$ we obtain the term
$\sum_{i=1}^{N}\lambda_{N}^{i,s}(\mathbf{x})$
which is not a function of $m_{N}^{\pm}(\mathbf{x})$ unless $f$ is linear.
@@ Appendix B Auxiliary computations
By comparison principle with the solution of the linear ODE (i.e. $B=0$) one can see that $y_{m}(t)>0$. Consider the function $\varphi(y):=A+By^{\alpha}-\frac{m}{2}y$. Note that $\varphi^{\prime}(y)<0$ for $y>y_{m}^{*}:=\left(\frac{2\alpha B}{m}\right)^{\frac{1}{1-\alpha}}$. Set now $Y_{m}:=\frac{2(A+B)}{m}$. Note that for $m$ sufficiently large, $Y_{m}>y_{m}^{*}$ and $\varphi(Y_{m})\leq 0$, so that $\varphi(y)<0$ for all $y\geq Y_{m}$. Let $\tau_{m}=\inf_{t>0}\{y_{m}(t)\leq\bar{Y}_{m}\}$. For $t<\tau_{m}$
@@
On the other hand, $y_{m}(t)\leq Y_{m}$ for $t\geq\tau_{m}$. Summing up, for all $t>0$ and $m$ sufficiently large,
@@
with $s(y)=\exp(-\int_{1}^{x}2\mu(y)/\sigma^{2}(y)dy)$. We have that $S(+\infty)=+\infty$ for any parameters choice, so the process does not explode, and $S(0)=$-$\infty$ if $c\geq a$, so in this case [35][Theorem 3.1] tells us that $P(T_{e}=+\infty)=1$, so there exist a (weak) unique (pathwise) solution to (84) and with probability one the process remains strictly positive for all times $t$ (the boundary is not attainable). If $0<c<a$, we have $S(0)\in{\mathbb{R}}$ and the process hits 0, i.e. $P(T_{0}<+\infty)=1$. We have $dX_{t}=cdt$, with $c>0$, when $X_{t}=0$, so the process is reflected at 0 and the explosion time is still infinity, so existence still holds. Pathwise uniqueness can be shown adapting [35][Theorem 3.2], using the fact that $bX^{2}<0$ to control $I_{1}$ in the proof. Note that this is consistent with the known results on CIR process.
To obtain a solution to the bi-dimensional equation, the second component must be given by
@@
To have weak existence and pathwise uniqueness, we check that $\sqrt{X_{\cdot}}\in M^{2}[0,T]$.

### v3

@@ 3.3 Convergence of the joint price-volatility process
The representations (70), (71) and (72) imply that the pair $(Z_{N},Y_{N})$ is not a Markov process. To see this we observe that (20), with $\lambda^{i,\pm}$ in place of $\lambda^{\pm}$ implies that the $0,+\infty)^{N}-$valued process $\mathbf{X}^{\pm}(t)=(X_{i}^{\pm}(t))_{i=1}^{N}$ is a Markov process whose infinitesimal generator $\mathcal{L}_{N}$ acts on a smooth $F:[0,+\infty)^{N}\rightarrow{\mathbb{R}}$ as follows: for $\mathbf{x}=(x_{1}^{\pm},\ldots,x_{N}^{\pm})$
@@
Note that if $(Z_{N}(t),Y_{N}(t))$ were a Markov process, then $m_{N}^{i,\pm}(\mathbf{X(t)})$ would also be a Markov process, since they are linked by an invertible linear transformation (except in the special case $\beta=\gamma=1$; but in this case $Z_{N}=0$ and the argument can be adapted by replacing the pair $m_{N}^{i,\pm}(\mathbf{X(t)})$ with the sum of the two components). This would imply that, for a smooth $\Psi:[0,+\infty)^{2}\rightarrow{\mathbb{R}}$, the generator $\mathcal{L}_{N}\Psi(m_{N}^{\pm}(\mathbf{x}))$ would be a function of $m_{N}^{\pm}(\mathbf{x})$. This is not the case. Indeed observe that, by permutation invariance, $T_{i}^{s}\Psi-\Psi$ does not depend on $i$, and therefore in $\mathcal{L}_{N}\Psi(m_{N}^{\pm}(\mathbf{x}))$ we obtain the term $\sum_{i=1}^{N}\lambda_{N}^{i,s}(\mathbf{x})$ which is not a function of $m_{N}^{\pm}(\mathbf{x})$ unless $f$ is linear.
@@ Appendix B Auxiliary computations
By comparison principle with the solution of the linear ODE (i.e. $B=0$) one can see that $y_{m}(t)>0$. Consider the function $\varphi(y):=A+By^{\alpha}-\frac{m}{2}y$. Note that $\varphi^{\prime}(y)<0$ for $y>y_{m}^{*}:=\left(\frac{2\alpha B}{m}\right)^{\frac{1}{1-\alpha}}$. Set now $Y_{m}:=\frac{2(A+B)}{m}$. Note that for $m$ sufficiently large, $Y_{m}>y_{m}^{*}$ and $\varphi(Y_{m})\leq 0$, so that $\varphi(y)<0$ for all $y\geq Y_{m}$. Let $\tau_{m}=\inf_{t>0}\{y_{m}(t)\leq{Y}_{m}\}$. For $t<\tau_{m}$
@@
On the other hand, $y_{m}(t)\leq Y_{m}$ for $t\geq\tau_{m}$, because $\dot{y}_{m}(t)\leq 0$ for $t\geq\tau_{m}$ . Summing up, for all $t>0$ and $m$ sufficiently large,
@@
with $s(y)=\exp(-\int_{1}^{x}2\mu(y)/\sigma^{2}(y)dy)$. We have that $S(+\infty)=+\infty$ for any parameters choice, so the process does not explode, and $S(0)=$-$\infty$ if $c\geq a$, so in this case [36][Theorem 3.1] tells us that $P(T_{e}=+\infty)=1$, so there exist a (weak) unique (pathwise) solution to (88) and with probability one the process remains strictly positive for all times $t$ (the boundary is not attainable). If $0<c<a$, we have $S(0)\in{\mathbb{R}}$ and the process hits 0, i.e. $P(T_{0}<+\infty)=1$. We have $dX_{t}=cdt$, with $c>0$, when $X_{t}=0$, so the process is reflected at 0 and the explosion time is still infinity, so existence still holds. Pathwise uniqueness can be shown adapting [36][Chapter IV - Theorem 3.2], using the fact that $bX^{2}<0$. To obtain a solution to the bi-dimensional equation, the second component must be given by
@@
To have weak existence and pathwise uniqueness, we check that